OpenStudy (anonymous):

3/8(7+2x)+3 greater then or equal too 18

OpenStudy (anonymous):

math is nothing but an illusion created by the government to kethe minds of homosapians at bay while they create a giant death starto blow up mars

OpenStudy (anonymous):

^keep

OpenStudy (anonymous):

multiply through, remove brackets

OpenStudy (anonymous):

3/8 (7+2x) +3 > or = to 18. Distribute. 3/8 (7) = 21/8 3/8 (2x) = 6x/8 So you have 21/8 + 6x/8 +3/1 is greater than or equal to 18.

OpenStudy (anonymous):

that dont make sense

OpenStudy (anonymous):

The only way you can add these and solve for X is to make 3 over 8 (a common denominator). So... multiply each term by 8. You get: 21 +6x +3 is greater than or equal to 0. 21+6x is greater than or equal to -3 Move the 21 to the other side. So: 6x is greater than or equal to -24 Divide by 6. X is greater than or equal to -4.

OpenStudy (anonymous):

Your grammar "don't make sense".

OpenStudy (anonymous):

Don't talk about my Gramma!

OpenStudy (anonymous):

xD